# BE Lyncis: An Extremely Eccentric Binary with the Nearest Known Black Hole

### v1

@@ III Conclusion
We have established BE Lyncis as an extraordinary binary system. The primary is a post-main sequence high-amplitude $\delta$ Scuti star with a mass of $1.55\,M_{\odot}$. Its unseen companion orbits with an unprecedented eccentricity of $e=0.9989^{+0.0008}_{-0.0021}$ (95% confidence >>0.9968), making this the most eccentric binary known to date. This extreme orbit, coupled with dynamical constraints that demand a companion mass $M_{2}\gtrsim 17.5\,M_{\odot}$ at an orbital inclination $i\lesssim 4.0^{\circ}$, unequivocally identifies the companion as a stellar-mass black hole. With a distance of approximately 250 pc, this black hole is the closest known to Earth.
The discovery of BE Lyn carries profound implications across several astrophysical frontiers. First, its record-breaking eccentricity serves as a direct observational constraint on the strength of natal kicks imparted during asymmetric supernova explosions, challenging models of black hole formation. Second, it is the first known binary comprising a pulsating star and a quiescent black hole, creating a unique laboratory for strong-gravity asteroseismology. Third, this work validates the pulsation-timing method—using the precise clocks of $\delta$ Scuti stars—as a novel, independent pathway to discover dormant black holes in wide binaries, thereby opening a new observational channel in the quest to unveil the hidden compact-object population.
Key questions regarding the exact black hole mass, the detailed evolutionary pathway leading to such an extreme configuration, and the system’s ultimate fate remain. Future multi-wavelength observations, astrometric monitoring, and theoretical modeling of supernova kicks in binary systems will be essential to fully understand the origins and implications of this remarkable cosmic benchmark.
Materials & Correspondence
Correspondence and requests for materials should be addressed to Jia-Shu Niu. (email: jsniu@sxu.edu.cn).
@@
{contribution}
J.S.N. and H.F.X. designed the research plan, performed the calculation, and wrote the manuscript. Y.Z. reviewed relevant literature, collected the TML, performed the frequency analysis, and wrote the initial version of Sec. 1 and part of Sec. 2.
Competing Interests
The authors declare that they have no competing financial interests.
Appendix A Methods
A.1 Photometric Data and Frequency Analysis
BE Lyn was observed by the Transiting Exoplanet Survey Satellite (TESS) in Sector 21 (January 2020) with a 120-second cadence. We retrieved the Presearch Data Conditioning Simple Aperture Photometry (PDCSAP) flux from the Mikulski Archive for Space Telescopes (MAST) and converted it to magnitudes using the TESS magnitude system (J. C. Smith et al., 2012; M. C. Stumpe et al., 2012, 2014; K. G. Stassun et al., 2019).
@@ A.1 Photometric Data and Frequency Analysis
where $m_{0}$ is the mean magnitude, and $a_{i}$, $f_{i}$, and $\phi_{i}$ are the amplitude, frequency, and phase of each component. Frequencies were accepted as significant if their signal-to-noise ratio (S/N) exceeded 5.6 (J.-S. Niu & H.-F. Xue, 2022; J.-S. Niu et al., 2023; J.-S. Niu & H.-F. Xue, 2024). After removing aliases, we identified 37 significant peaks, consisting of 7 independent modes and 30 harmonics or linear combinations (Table A1).
A.2 Compilation and Processing of Times of Maximum Light
A.3 Orbital Modeling via MCMC
The OCO-C residuals were modeled with a function combining a quadratic term (intrinsic period variation) and a light-travel time effect (LTTE) term for orbital motion (H.-F. Xue & J.-S. Niu, 2020):
$C=\mathrm{BJD}_{0}+P_{0}E+\frac{1}{2}\beta E^{2}+A\left[\sqrt{1-e^{2}}\sin{\phi}\cos{\omega}+\cos{\phi}\sin{\omega}\right],$ (A2)
A.4 Asteroseismic Modeling of the Primary Star
The optimal model was selected by minimizing $\chi^{2}$ across three observables: the fundamental frequency ($f_{0}$), first-overtone frequency ($f_{1}$), and the intrinsic period variation rate $(1/P_{0})(dP_{0}/dt)$. This methodology follows established asteroseismic analyses of HADS (J.-S. Niu et al., 2017; H.-F. Xue et al., 2018, 2022). The best-fit model parameters are given in Table A2.
A.5 Dynamical Constraints on the Companion
The companion’s nature was inferred from the mass function,
$f(M_{2})=\frac{(M_{2}\sin i)^{3}}{(M_{1}+M_{2})^{2}}=\frac{4\pi^{2}}{GP_{\mathrm{orb}}^{2}}(a\sin i)^{3},$ (A3)
combined with the constraints that the periastron separation $d_{\mathrm{peri}}=a(1-e)(M_{1}+M_{2})/M_{2}$ must exceed the primary’s radius $R_{1}$, and the Roche lobe radius $R_{1,\mathrm{R}}$ must exceed the primary’s radius $R_{1}$. The Roche lobe radius was calculated using the approximation of P. P. Eggleton (1983). This yields an upper limit on the inclination ($i\lesssim 4.0^{\circ}$) and a corresponding lower limit on the companion mass ($M_{2}\gtrsim 17.5M_{\odot}$), as shown in Figure 3.
Appendix B Additional Figures
Figure A1: Light curve of BE Lyn from TESS Sector 21. The lower panel shows a one-day zoom for clarity.
Figure A2: Frequency spectrum of BE Lyn. The main panel (0100c/d100\ $\mathrm${c/d}) marks the fundamental ($f_{0}$, red) and first-overtone ($f_{1}$, green) modes. The inset highlights the low-amplitude, non-radial mode region ($\sim 9-17\ \mathrm{c/d}$) with a linear y-axis.
Appendix C Additional Tables
Table A1: Frequency solution for BE Lyn from TESS Sector 21.
@@ Appendix C Additional Tables
F13 83.4470 0.0006 0.47 0.01 31.8 $8f_{0}$
F14 7.440 0.001 0.37 0.02 21.3 $2f_{0}-f_{1}$
F15 44.7150 0.0008 0.29 0.01 24.4 $3f_{0}+f_{1}$
F16 93.8765 0.0009 0.26 0.01 23.8 $9f_{0}$
F19 55.147 0.001 0.16 0.01 17.7 $4f_{0}+f_{1}$
F20 9.701 0.003 0.15 0.02 7.2 $f_{2}$
F22 104.309 0.002 0.14 0.01 14.0 $10f_{0}$
F24 17.869 0.002 0.11 0.01 10.5 $3f_{0}-f_{1}$
F25 65.579 0.002 0.10 0.01 10.4 $5f_{0}+f_{1}$
F26 10.684 0.003 0.10 0.02 6.0 $f_{3}$
F27 9.377 0.003 0.09 0.01 6.0 $f_{4}$
F28 114.740 0.002 0.076 0.008 9.8 $11f_{0}$
F29 14.009 0.003 0.08 0.01 6.4 $f_{5}$
F30 19.810 0.003 0.08 0.01 6.2 $f_{0}+f_{4}$
F31 20.133 0.003 0.07 0.01 5.9 $f_{0}+f_{2}$
F32 21.112 0.004 0.070 0.01 5.7 $f_{0}+2f_{3}$
F33 76.006 0.003 0.067 0.009 7.5 $6f_{0}+f_{1}$
F34 16.336 0.003 0.07 0.01 6.8 $f_{6}$
F35 28.299 0.003 0.058 0.009 6.1 $4f_{0}-f_{1}$
F36 38.729 0.003 0.052 0.008 6.1 $5f_{0}-f_{1}$
F37 86.438 0.004 0.045 0.008 5.8 $7f_{0}+f_{1}$
Table A2: Parameters of the best-fit seismic model for BE Lyn’s primary star.
@@
$f_{0}$ (c/d) 10.4288 $10.4309\pm 0.0005$
$f_{1}$ (c/d) 13.4243 $13.4226\pm 0.0005$
$(1/P_{0})(dP_{0}/dt)$ (yr-1) $1.34\times 10^{-8}$ $(1.02^{+0.08}_{-0.08})\times 10^{-8}$
$\log T_{\mathrm{eff}}$ 3.902
$\log(L/L_{\odot})$ 1.326
$\log g$ 3.865
$\log(R/R_{\odot})$ 0.382
Age (Gyr) 1.59
Mass ($M_{\odot}$) 1.55
@@
Table A3: Times of maximum light of BE Lyn. The numbers in the fourth, eighth, and twelfth columns indicate the different data sources: (1) T. Oja (1987); (2) E. Rodríguez et al. (1990); (3) Y. Liu et al. (1991); (4) E. Wunder et al. (1992); (5) Q. Tang et al. (1992); (6) Z. Liu & S. Jiang (1994); (7) L. L. Kiss & K. Szatmary (1995); (8) A. Derekas et al. (2003); (9) J. N. Fu & S. Y. Jiang (2005); (10) R. Szakats et al. (2008); (11) J. Hubscher (2005);(12) J. Hubscher et al. (2005); (13) J. Hubscher et al. (2006); (14) J. Hübscher et al. (2005); (15) G. Klingenberg et al. (2006); (16) C. Boonyarak et al. (2011); (17) J. Hubscher (2007); (18) J. Hubscher et al. (2009); (19) AAVSO; (20) J. Hubscher et al. (2013); (21) J. H. Peña et al. (2015); (22) J. Hubscher (2014); (23) J. Hübscher (2017); (24) TESS; (25) L. Pagel (2022); (26) L. Pagel (2021); (27) L. Pagel (2023).

### v2

@@ III Conclusion
We have established BE Lyncis as an extraordinary binary system. The primary is a post-main-sequence, high-amplitude $\delta$ Scuti star with a mass of $1.61\,M_{\odot}$. Its unseen companion follows an orbit with an eccentricity of $e=0.9989^{+0.0008}_{-0.0021}$ (95% confidence >0.9968>0.9968), the most extreme reliably measured in any binary system to date. Dynamical constraints, rooted in orbital stability, limit the inclination to $i\lesssim 10.1^{\circ}$ and demand a companion mass $M_{2}\gtrsim 2.5\,M_{\odot}$, placing it squarely spanning the mass gap between neutron stars and black holes.
The discovery of BE Lyn carries profound implications across several astrophysical frontiers. First, its record-breaking eccentricity provides a stringent observational benchmark for models of natal kicks in asymmetric supernovae, challenging our understanding of compact-object formation and evolution. Second, it represents the first known binary system to pair a pulsating star with a quiescent compact object spanning the mass gap, opening the door to strong-gravity asteroseismology—a novel probe of spacetime around massive companions. Third, our work validates the pulsation-timing method, which harnesses the precise clocks of $\delta$ Scuti stars, as an independent and powerful pathway to discover dormant compact objects in wide binaries. This technique complements existing radial-velocity and astrometric surveys and offers a new observational window into the hidden population of neutron stars and black holes.
Fundamental questions remain. The exact mass and nature of the companion—whether the most massive neutron star known or the nearest black hole to Earth—await definitive measurement. The evolutionary pathway that produced such an extreme configuration, and the ultimate fate of the system, are yet to be understood. Future multi-wavelength observations, astrometric monitoring (e.g., with Gaia and future missions), and theoretical modeling of supernova kicks and binary evolution will be essential to fully unravel the origins and implications of this remarkable cosmic benchmark.
Acknowledgments
@@
Appendix A Photometric Data and Frequency Analysis
BE Lyn was observed by the Transiting Exoplanet Survey Satellite (TESS) in Sector 21 (January 2020) with a 120-second cadence. We retrieved the Presearch Data Conditioning Simple Aperture Photometry (PDCSAP) flux from the Mikulski Archive for Space Telescopes (MAST) and converted it to magnitudes using the TESS magnitude system (J. C. Smith et al., 2012; M. C. Stumpe et al., 2012, 2014; K. G. Stassun et al., 2019) (see in Figure 4).
Figure 4: Light curve of BE Lyn from TESS Sector 21. The lower panel shows a one-day zoom for clarity.
@@ Appendix A Photometric Data and Frequency Analysis
where $m_{0}$ is the mean magnitude, and $a_{i}$, $f_{i}$, and $\phi_{i}$ are the amplitude, frequency, and phase of each component. Frequencies were accepted as significant if their signal-to-noise ratio (S/N) exceeded 5.6 (W. Zong et al., 2016, 2018). After removing aliases, we identified 32 significant peaks, consisting of 7 independent modes and 25 harmonics or linear combinations (see in Figure 5 and Table 2).
Figure 5: Frequency spectrum of BE Lyn. The main panel (0100c/d100\ $\mathrm${c/d}) marks the fundamental ($f_{0}$, red) and first-overtone ($f_{1}$, green) modes. The inset highlights the low-amplitude, non-radial mode region ($\sim 9-17\ \mathrm{c/d}$) with a linear y-axis.
Table 2: Significant Frequencies Detected in BE Lyn from TESS Sector 21.
@@
F12 83.4470 0.0006 0.47 0.01 31.8 $8f_{0}$
F13 7.440 0.001 0.37 0.02 21.3 $2f_{0}-f_{1}$
F14 44.7150 0.0008 0.29 0.01 24.4 $3f_{0}+f_{1}$
F15 93.8765 0.0009 0.26 0.01 23.8 $9f_{0}$
F16 55.147 0.001 0.16 0.01 17.7 $4f_{0}+f_{1}$
F17 9.701 0.003 0.15 0.02 7.2 $f_{2}$
F18 104.309 0.002 0.14 0.01 14.0 $10f_{0}$
F19 17.869 0.002 0.11 0.01 10.5 $3f_{0}-f_{1}$
F20 65.579 0.002 0.10 0.01 10.4 $5f_{0}+f_{1}$
F21 10.684 0.003 0.10 0.02 6.0 $f_{3}$
F22 9.377 0.003 0.09 0.01 6.0 $f_{4}$
F23 114.740 0.002 0.076 0.008 9.8 $11f_{0}$
F24 14.009 0.003 0.08 0.01 6.4 $f_{5}$
F25 19.810 0.003 0.08 0.01 6.2 $f_{0}+f_{4}$
F26 20.133 0.003 0.07 0.01 5.9 $f_{0}+f_{2}$
F27 21.112 0.004 0.07 0.01 5.7 $f_{0}+2f_{3}$
F28 76.006 0.003 0.067 0.009 7.5 $6f_{0}+f_{1}$
F29 16.336 0.003 0.07 0.01 6.8 $f_{6}$
F30 28.299 0.003 0.058 0.009 6.1 $4f_{0}-f_{1}$
F31 38.729 0.003 0.052 0.008 6.1 $5f_{0}-f_{1}$
F32 86.438 0.004 0.045 0.008 5.8 $7f_{0}+f_{1}$
Appendix B Compilation and Processing of Times of Maximum Light
Appendix C Orbital Modeling via MCMC
The OCO-C residuals were modeled with a function combining a quadratic term (intrinsic period variation) and a light-travel time effect (LTTE) term for orbital motion (M. Paparo et al., 1988; H.-F. Xue & J.-S. Niu, 2020):
[…]
Appendix D Asteroseismic Modeling of the Primary Star
The optimal model was selected by minimizing $\chi^{2}$ across two observables: the fundamental frequency ($f_{0}$) and first-overtone frequency ($f_{1}$). This methodology follows established asteroseismic analyses of HADS (J.-S. Niu et al., 2017; H.-F. Xue et al., 2018, 2022). The best-fit model parameters are given in Table 3.
Table 3: Parameters of the best-fit seismic model for BE Lyn’s primary star.
@@ Appendix D Asteroseismic Modeling of the Primary Star
$f_{0}$ (c/d) 10.4310 $10.4309\pm 0.0005$
$f_{1}$ (c/d) 13.4226 $13.4226\pm 0.0005$
$(1/P_{0})(dP_{0}/dt)$ (yr-1) $8.36\times 10^{-8}$ $(1.02^{+0.08}_{-0.08})\times 10^{-8}$
$\log T_{\mathrm{eff}}$ 3.915
$\log(L/L_{\odot})$ 1.391
$\log g$ 3.870
$\log(R/R_{\odot})$ 0.387
Age (Gyr) 1.44
Mass ($M_{\odot}$) 1.61
Appendix E Dynamical Constraints on the Companion
The companion’s nature was inferred from the mass function,
$f(M_{2})=\frac{(M_{2}\sin i)^{3}}{(M_{1}+M_{2})^{2}}=\frac{4\pi^{2}}{GP_{\mathrm{orb}}^{2}}(a\sin i)^{3},$ (E1)
combined with the constraints that the periastron separation $d_{\mathrm{peri}}=a(1-e)(M_{1}+M_{2})/M_{2}$ must exceed the primary’s radius $R_{1}$. This yields an upper limit on the inclination ($i\lesssim 10.1^{\circ}$) and a corresponding lower limit on the companion mass ($M_{2}\gtrsim 2.5M_{\odot}$), as shown in Figure 3.
Appendix F Additional Tables
All the TML collected and used in this work are listed in Table 4.
@@ Appendix F Additional Tables
Table 4: Times of maximum light of BE Lyn. The numbers in the fourth, eighth, and twelfth columns indicate the different data sources: (1) T. Oja (1987); (2) E. Rodríguez et al. (1990); (3) Y. Liu et al. (1991); (4) E. Wunder et al. (1992); (5) Q. Tang et al. (1992); (6) Z. Liu & S. Jiang (1994); (7) L. L. Kiss & K. Szatmary (1995); (8) A. Derekas et al. (2003); (9) J. N. Fu & S. Y. Jiang (2005); (10) R. Szakats et al. (2008); (11) J. Hubscher (2005);(12) J. Hubscher et al. (2005); (13) J. Hubscher et al. (2006); (14) J. Hübscher et al. (2005); (15) G. Klingenberg et al. (2006); (16) C. Boonyarak et al. (2011); (17) J. Hubscher (2007); (18) J. Hubscher et al. (2009); (19) AAVSO; (20) J. Hubscher et al. (2013); (21) J. H. Peña et al. (2015); (22) J. Hubscher (2014); (23) J. Hübscher (2017); (24) TESS; (25) L. Pagel (2022); (26) L. Pagel (2021); (27) L. Pagel (2023).